Ching leaves San Francisco and travels toward Los Angeles at 50 mi/h. An hour later, Phil leaves Los Angeles and travels toward San Francisco at 60 mi/h. If the two cities are 380 miles apart, how many hours will it take for Ching to meet Phil?
Answer by checkley75(3666)

50x+60(x-1)=380
50x+60x-60=380
110x=+380+60
110x=440
x=440/110
x=4 hours for Phil meet up with Ching.
proof
50*4+60*3=380
200+180=380
380=380